OpenStudy (anonymous):

Expressed this in fraction. .090909090909........

OpenStudy (anonymous):

this is: \[\frac{9}{100}+\frac{9}{10000}+\frac{9}{1000000}\cdots\] First term is 9/100, r = 1/100 Use the formula for the sum of an infinite series.

OpenStudy (anonymous):

so, what is the final asnwer?

OpenStudy (anonymous):

I doubt alfers knows Calc II, joe. The more basic way: x = 0.09... 100x = 9.09... - x = 0.09... ------------ 99x = 9 x = 9/99 = 1/11.

OpenStudy (anonymous):

yeah! thank u so much. i remember now. =] thank u!!

OpenStudy (anonymous):

1/11